Part C Broa (aq) N2H4 (g) Br (ag) N2 (g) Express your answer as a chemical equation. Identify all of the ANSWER
Answer #1

Split

BrO3- = Br-

N2H4 = N2

balance O

BrO3- = Br- + 3H2O

N2H4 = N2

balance H

6H+ + BrO3- = Br- + 3H2O

N2H4 = N2 + 4H+

balance charge

6e- + 6H+ + BrO3- = Br- + 3H2O

N2H4 = N2 + 4H+ + 4e-

balance e

12e- + 12H+ + 2BrO3- = 2Br- + 6H2O

3N2H4 = 3N2 + 12H+ + 12e-

add all

3N2H4 + 12e- + 12H+ + 2BrO3- = 2Br- + 6H2O +  3N2 + 12H+ + 12e-

cancel e-

3N2H4 + 2BrO3- = 2Br- + 6H2O +  3N2